OpenStudy (anonymous):

function symbol= 3x + 2 and g(x) = 6x – 7. Find function symbol – g(x) 3x – 9 –3x + 9 –3x – 5 2x – 1

OpenStudy (anonymous):

let function symbol=y(x) so, y(x)=3x+2 and g(x)=6x-7 so y(x)-g(x)=(3x+2)-(6x-7) =3x+2-6x+7 =3x-6x+2+7 =-3x+9